# Intermittent time series forecasting with Gaussian Processes and Tweedie likelihood

### v4

@@ Abstract
Intermittent time series , Gaussian Processes , Tweedie distribution , Probabilistic forecasting
@@
[1]organization=SUPSI, Istituto Dalle Molle di Studi sull’Intelligenza Artificiale (IDSIA), city=Lugano,country=Switzerland
@@ 1 Introduction
Intermittent time series characterize a large percentage of the inventory items. Traditional forecasting methods for intermittent demand (Croston, 1972; Syntetos and Boylan, 2005; Nikolopoulos et al., 2011) provide point forecasts only. However, planning the inventory levels requires probability distributions from which to extract the relevant quantiles (Boylan and Syntetos, 2021; Kolassa, 2016).
@@
Probabilistic models for intermittent demand, such as Hyndman et al. (2008, pp.281-283), Snyder et al. (2012), Sbrana (2023), and Svetunkov and Boylan (2023), return the forecast distribution $p(y_{T+1}\mid y_{1:T})$, where $y_{1:T}$ are the observations available up to time $T$ and $y_{T+1}$ is the predicted value at time $T+1$. Such models predict the distribution of the demand size (the positive values of the time series) and the probability of occurrence, i.e., the binary variable indicating whether there will be positive demand.
@@
Most probabilistic methods in the literature use one or more stochastic processes to control how the parameters of the forecast distribution change over time. The realisation of such processes is a function of time, and is latent since cannot be observed. For instance in the local level model (Hyndman et al., 2008, Chap. 3.4), the latent function parametrizes the mean of the forecast distribution.
@@
Yet, the above models do not account for the uncertainty about the value of the latent variables, which might be substantial. Indeed, incorporating the uncertainty on the parameters of the forecast distribution leads to more reliable forecasts (Prak and Teunter, 2019); this is a typical advantage of the Bayesian approach (Gelman et al., 2020, Chap. 9).
@@
We propose to model the latent variable with a Gaussian Process (GP), a Bayesian non-parametric model which quantifies the uncertainty of the latent variable and propagates it to the forecast distribution. GPs have been previously applied to smooth time series (Roberts et al., 2013; Corani et al., 2021), but not yet to intermittent time series. Dealing with smooth time series, the GP latent function is generally coupled with a Gaussian forecast distribution. This is however not suitable for intermittent time series, which require a distribution defined over positive values and with a mass in zero. Also the Poisson is not a suitable forecast distribution for intermittent demand because of its short tails, which do not cover demand spikes.
@@
A suitable distribution is, instead, the the negative binomial (Harvey and Fernandes, 1989; Snyder et al., 2012; Kolassa, 2016). Indeed, our first model (NegBinGP) couples the GP with a negative binomial forecast distribution. The negative binomial is however unimodal; this property might be restrictive. The previously mentioned probabilistic models for intermittent time series have a bimodal distribution, with a point mass in zero and a distribution over positive values.
In our second model (TweedieGP) the GP parametrizes the mean of a Tweedie (Dunn and Smyth, 2005) distribution, which is a bimodal distribution with a mode in zero and a long right tail.
@@
TweedieGP is the first probabilistic model for intermittent time series based on a fully parameterized Tweedie distribution. We also show (Sec. 3.2.1) that the Tweedie loss, often used for training point forecast models on intermittent time series (Jeon and Seong, 2022; Januschowski et al., 2022), is obtained by severely approximating the Tweedie distribution.
@@
We perform experiments on about 40’000 supply chain time series. Both GPs generally provide better probabilistic forecasts than the competitors. In particular, TweedieGP often outperforms the competitors on the highest quantiles, the most difficult to estimate. This might be due to additional flexibility of the Tweedie distribution compared to the negative binomial. Thanks to variational methods (Hensman et al., 2015), the training times of GPs are in line with those of other local models for intermittent time series.
Our implementation is based on GPyTorch (Gardner et al., 2018). Upon acceptance we will publicly release the implementation of our models and of the Tweedie distribution.
@@ 2 Literature on probabilistic models for intermittent time series
The simplest approach for forecasting intermittent time series is a static model. This is appropriate (Snyder et al., 2012) for time series containing a very large amount of zeros. The static model can be the empirical distribution (Boylan and Syntetos, 2021, Sec 13.2) or a fitted distribution such as the negative binomial (Kolassa, 2016).
@@
$d_{t}$ :=$:=\begin{cases}y_{t}\ &\text{if}\ y_{t}>0\\ \text{undefined}&\text{otherwise}\end{cases}\quad{\color[rgb]{0.,0.,0.}.}$
@@
Thus the values of a time series can be written as $y_{t}=o_{t}\cdot d_{t}$.
@@
Several models are based on this decomposition. Croston’s method (Croston, 1972), predicts demand sizes and demand intervals with two independent exponential smoothing processes. The TSB method (Teunter et al., 2011) improves by using the second exponential smoothing to predict the occurrence rather than the demand interval. Both models return point forecasts only. Their probabilistic counterparts have been developed by Hyndman et al. (2008) and Snyder et al. (2012) respectively. The forecast distribution of both models is bimodal, and it is given by a mixture of a Bernoulli and a shifted Poisson distribution.
@@
Snyder et al. (2012) propose a further model with a single latent exponential smoothing variable that controls the mean of a negative binomial distribution; in this case the forecast distribution is unimodal and overdispersed.
@@
Sbrana (2023) adopts a latent process which can be equal to 0 with constant probability. The resulting forecast distribution is a mixture of a mass in zero and a truncated Gaussian distribution.
@@
iETS (Svetunkov and Boylan, 2023) is possibly the most sophisticated model in this class. It assumes again the independence of occurrence and demand size and it models the latent variables via exponential smoothing. The latent demand size variable is modelled as a multiplicative exponential smoothing process. iETS considers different models of occurrence, which are based on one or two processes; they cover cases such as demand building up, demand obsolescence, etc. The best occurrence model is chosen via AICc. The forecast distribution is the mixture of a Bernoulli and a Gamma distribution.
@@
The limit of the above models methods is that they do not account for the uncertainty of the latent processes. Bayesian models (Yelland, 2009; Chapados, 2014; Babai et al., 2021) overcome this limitation, but there is no public implementation of them, and they can be computationally intensive.
@@
WSS (Willemain et al., 2004) is a non-parametric approach based on the independence between occurrence and demand, which models the occurrence using a two-states Markov chain. When the simulated occurrence is positive, it samples the demand size from past values via bootstrapping with jittering. The forecast distribution is a mixture of a mass in zero and an integer-valued, non-parametric distribution. A limit of WSS is that it does not model the dynamic of the demand size.
@@
It is thus common to independently model occurrence and demand size, which results in a bimodal forecast distribution with a peak mass in zero to model the absence of demand, and a positive mode to describe demand size. However, on sparse time series, the estimate of the demand size is only based on few observations, which can result in large uncertainty.
@@
ADIDA (Nikolopoulos et al., 2011) temporally aggregates the intermittent time series. The aggregated time series are simpler to forecast as they contain fewer zeros. The predictions on aggregated time series, made with the naïve method, are then disaggregated to the original time buckets. ADIDA can be coupled with bootstrapping or conformal inference to produce probabilistic forecasts.
@@
Combinations of point forecasts (Petropoulos and Kourentzes, 2015) and probabilistic forecasts (Wang et al., 2025) have been used to stabilize the performance of above methods. The computational cost is higher, but model selection is not required.
@@ 3 Gaussian Processes
The models reviewed above often assume a specific parametric form for their latent variables. Moreover, they often estimate it with a point forecast. Gaussian processes (GP, Rasmussen and Williams, 2006) model the latent function in a non-parametric way, thus avoiding strong assumptions. Furthermore, a Gaussian process is a Bayesian method: parameters are treated as random variables, to which a prior distribution is assigned. Uncertainty is then updated based on observed data using Bayes’ theorem, resulting in a posterior distribution. The distribution of the data given the parameters is known as the likelihood (Gelman et al., 2020, Chap. 9).
@@ A GP provides a prior distribution on the space of functions $f:\mathbb{R}_{+}\rightarrow\mathbb{R}$:
The kernel $k$ is a positive definite function (Chap. 4, Rasmussen and Williams, 2006) and encodes the covariance between the values of $f$ at any two instants in time. We assume $f$ to be a smooth function of time. This assumption is encoded by the Radial Basis Function (RBF) kernel:
@@
By defining a prior in the space of functions, GPs do not need regularly spaced observations, therefore we consider a training set of $T$ couples $(t_{1},y_{1}),\ldots,(t_{T},y_{T})$, where, for each $i$, $t_{i}$ is the time stamp and $y_{i}$ is the observation. We aim to generate forecast for time inputs $T+1,\dots,T+h$, where $h$ is known as forecast horizon. We denote by $p(\cdot)$ the density functions. The GP provides a prior on the latent vector $\mathbf{f}_{1:T}:=(f(t_{1}),\ldots,f(t_{T}))^{\top}$ which, given a set of time instants $t_{1},\ldots,t_{T}$, encodes the value of the latent function at those times. The prior is a multivariate Gaussian distribution:
@@
Samples of vectors $\mathbf{f}_{1:T}$ are shown in Fig. 1 (left). Credible intervals display the spread of the prior distribution: at time $t$, the $q$-level credible interval is the interval that contains a proportion $q$ of values of $f(t)$. A priori, the function has the same mean and variance at any time; hence the credible intervals are flat (Fig. 1, right).
@@
The latent function $f$, defined by the GP prior in eq. (1), is a time-varying parameter related to the mean of the likelihood function, specified here as a negative binomial, (Sec. 3.1) or a Tweedie distribution (Sec. 3.2). For this reason, we require a positive latent function and we ensure this by passing $f$ through the softplus function, defined as $\mathrm{softplus}(x):=\log(1+e^{x})$. The softplus maps negative values of $x$ to small positive values, while it is close to the identity function for $x>2$. A priori, the credible intervals of $\mathrm{softplus}(f)$ are identical for all time instants, non-negative and asymmetric (Fig. 2, left).
@@
where $f_{i}:=f(t_{i})$, $p_{\mathrm{lik}}(\cdot\mid\mathrm{softplus}(f_{i}),\boldsymbol{\theta}_{\mathrm{lik}})$ is either the negative binomial or the Tweedie distribution and $\boldsymbol{\theta}_{\mathrm{lik}}$ are their hyper-parameters, which we discuss in Secs. 3.1 and 3.2.
@@
The posterior in eq. (4) is available in analytical form only for the Gaussian likelihood. Since our likelihoods are non-Gaussian, we approximate $p(\mathbf{f}_{1:T}~{}\mid~{}\mathbf{y}_{1:T})$ with a variational inducing point approximation (Hensman et al., 2013) constituted by a Gaussian density, denoted by $q(\mathbf{f}_{1:T}\mid\mathbf{y}_{1:T})$, with tunable mean and covariance parameters. We follow Hensman et al. (2015) to optimize the model parameters as described in A. In what follows we denote the fitted variational approximation by $q(\mathbf{f}_{1:T}\mid\mathbf{y}_{1:T})$.
@@
The distribution of the latent function at future times, shown in eq. (5), is obtained marginalizing a multivariate Gaussian distribution. Given a likelihood expressed as in eq. (3), $h$-steps ahead forecasts are obtained without autoregressive sampling. We now complete the GP model by specifying the likelihood function.
@@ 3.1 Negative binomial likelihood
The negative binomial is a suitable distribution for intermittent demand (Snyder et al., 2012; Kolassa, 2016). It addresses the limitations of the Poisson distribution since it can have both a large probability mass in zero and long tails.
We denote by $\mathrm{NegBin}(y;r,\text{p})$ the density of a negative binomial distribution; $r$ is the number of successes and p denotes, with a slight abuse of notation, the success probability. The mean and the variance of the distribution linearly increase with $r$; p controls the overdispersion. We model $r$ through the GP, by setting $r=\text{softplus}(f)$. Instead, we treat p as a learnable hyper-parameter ($\boldsymbol{\theta}_{\mathrm{lik}}=\{\text{p}\}$) which we keep fixed for all the data points of the same time series.
@@
$_{\mathrm{lik}}\left(y_{i}\mid\mathrm{softplus}(f_{i}),\boldsymbol{\theta}_{\mathrm{lik}}\right)=\mathrm{NegBin}(y_{i};\mathrm{softplus}(f_{i}),\text{p}).$ p ) .

### v5

# Lugano, Switzerland

@@ Abstract
Intermittent time series , Gaussian Processes , Tweedie distribution , Probabilistic forecasting , Quantile loss , Machine learning , Bayesian methods , Supply chain
@@ 1 Introduction
Intermittent time series characterize a large percentage of inventory items. Traditional forecasting methods for intermittent demand (Croston_1972; Syntetos_Boylan_2005; Nikolopoulos_Syntetos_Boylan_Petropoulos_Assimakopoulos_2011) provide point forecasts only. However, planning inventory levels requires probability distributions from which to extract the relevant quantiles (boylan2021intermittent; Kolassa_2016).
@@
Probabilistic models for intermittent demand, such as Hyndman_2008, snyder2012forecasting, sbrana2023modelling, and Svetunkov_Boylan_2023, return the forecast distribution $p(y_{T+1}\mid y_{1:T})$, where $y_{1:T}$ are the observations available up to time $T$ and $y_{T+1}$ is the predicted value at time $T+1$. Such models predict the distribution of the demand size (the positive values of the time series) and the probability of occurrence, i.e., the binary variable indicating whether there will be positive demand.
@@
Most probabilistic methods in the literature use one or more stochastic processes to control how the parameters of the forecast distribution change over time. Such stochastic processes are a function of time and their realisations are not observed, thus they are called latent. For instance in the local level model (Hyndman_2008, Chap. 3.4), the latent function parametrizes the mean of the forecast distribution.
@@
Yet, the above models do not account for the uncertainty about the value of the latent variables, which might be substantial. Indeed, incorporating the uncertainty on the parameters of the forecast distribution leads to more reliable forecasts (prak2019general); this is a typical advantage of the Bayesian approach (Gelman_Hill_Vehtari_2020, Chap. 9).
@@
We propose to model the latent variable with a Gaussian Process (GP), a Bayesian non-parametric model which quantifies the uncertainty of the latent variable and propagates it to the forecast distribution. GPs have been previously applied to smooth time series (Roberts_Osborne_Ebden_Reece_Gibson_Aigrain_2013; Corani_Benavoli_Zaffalon_2021), but not yet to intermittent time series. Dealing with smooth time series, the GP latent function is generally coupled with a Gaussian forecast distribution. However, this is not suitable for intermittent time series, which require a distribution defined over positive values and with a mass in zero. Also the Poisson distribution is not a suitable forecast distribution for intermittent demand because of its short tails, which do not cover demand spikes.
@@
A suitable distribution is, instead, the negative binomial (Harvey_Fernandes_1989; snyder2012forecasting; Kolassa_2016). Indeed, our first model (NegBinGP) couples the GP with a negative binomial forecast distribution. However, this distribution is unimodal; this property might be restrictive. Several probabilistic models for intermittent time series have a bimodal distribution, with a point mass in zero and a distribution over positive values.
@@
In our second model (TweedieGP) the GP parametrizes the mean of a Tweedie (Dunn_Smyth_2005) distribution, which is a bimodal distribution with a mode in zero and a long right tail. TweedieGP is the first probabilistic model for intermittent time series based on a fully parameterized Tweedie distribution. We also show (Sec. 3.2.1) that the Tweedie loss, often used for training point forecast models on intermittent time series (jeon2022robust; januschowski2022forecasting), is obtained by severely approximating the Tweedie distribution.
@@
We perform experiments on about 40’000 supply chain time series. Both GPs generally provide better probabilistic forecasts than the competitors. In particular, TweedieGP often outperforms the competitors on the highest quantiles, the most difficult to estimate. This might be due to the additional flexibility of the Tweedie distribution compared to the negative binomial. Thanks to variational methods (Hensman_Matthews_Ghahramani_2015), the training times of GPs are in line with those of other local models for intermittent time series.
Our implementation is based on GPyTorch (Gardner_Pleiss_Weinberger_Bindel_Wilson_2018). Our models and the Torch implementation of the Tweedie distribution are available at the GitHub page of the project [^3] https://github.com/StefanoDamato/TweedieGP.
@@ 2 Literature on probabilistic models for intermittent time series
The simplest approach for forecasting intermittent time series is a static model. This is appropriate (snyder2012forecasting) for time series containing a very large amount of zeros. The static model can be the empirical distribution (boylan2021intermittent, Sec 13.2) or a fitted distribution such as the negative binomial (Kolassa_2016).
@@
$d_{t}$ $:=\begin{cases}y_{t}\ &\text{if}\ y_{t}>0\\ \text{undefined}&\text{otherwise}\end{cases}\quad{\color[rgb]{0.,0.,0.},}$
@@
respectively. Thus the observed values of a time series can be written as $y_{t}=o_{t}\cdot d_{t}$.
@@
Several models are based on this decomposition. Croston’s method (Croston_1972), predicts demand sizes and demand intervals with two independent exponential smoothing processes. The TSB method (Teunter_Syntetos_Babai_2011) improves by using the second exponential smoothing to predict the occurrence rather than the demand interval. Both models return point forecasts only. Their probabilistic counterparts have been developed by Hyndman_2008 and snyder2012forecasting respectively. The forecast distribution of both models is bimodal, and it is given by a mixture of a Bernoulli and a shifted Poisson distribution.
@@
snyder2012forecasting propose a further model with a single latent exponential smoothing variable that controls the mean of a negative binomial distribution; in this case the forecast distribution is unimodal and overdispersed.
@@
sbrana2023modelling adopts a latent process which can be equal to zero with constant probability. The resulting forecast distribution is a mixture of a mass in zero and a truncated Gaussian distribution.
@@
iETS (Svetunkov_Boylan_2023) is possibly the most sophisticated model in this class. It assumes again the independence of occurrence and demand size. The latent demand size variable is modelled as a multiplicative exponential smoothing process. iETS considers different models of occurrence, which are based on one or two processes; they cover cases such as demand building up, demand obsolescence, etc. The best occurrence model is chosen via AICc. The forecast distribution is the mixture of a Bernoulli and a Gamma distribution.
@@
The limit of the above models methods is that they do not account for the uncertainty of the latent processes. Bayesian models (yelland2009bayesian; chapados2014effective; Babai_Chen_Syntetos_Lengu_2021) overcome this limitation, but there is no public implementation of them, and they can be computationally intensive.
@@
WSS (Willemain_Smart_Schwarz_2004) is a non-parametric approach based on the independence between occurrence and demand, which models the occurrence using a two-states Markov chain. When the simulated occurrence is positive, it samples the demand size from past values via bootstrapping with jittering. The forecast distribution is a mixture of a mass in zero and an integer-valued, non-parametric distribution. A limit of WSS is that it does not model the dynamic of the demand size.
@@
It is thus common to independently model occurrence and demand size, leading to a bimodal forecast distribution with a peak mass in zero to predict the absence of demand, and a positive mode to describe demand size. However, on sparse time series, the estimate of the demand size is only based on few observations, which can result in large uncertainty.
@@
ADIDA (Nikolopoulos_Syntetos_Boylan_Petropoulos_Assimakopoulos_2011) temporally aggregates the intermittent time series. The aggregated time series are simpler to forecast as they contain fewer zeros. The predictions on these series, made with the naïve method, are then disaggregated to the original time buckets. ADIDA can be coupled with bootstrapping or conformal inference to produce probabilistic forecasts.
@@
Combinations of point (Petropoulos_Kourentzes_2015) and probabilistic forecasts (Wang_Kang_Petropoulos_2024) have been used to stabilize the performance of above methods. The computational cost is higher, but model selection is not required.
@@ 3 Gaussian Processes
The models reviewed above often assume a specific parametric form for their latent variables. Moreover, they often estimate it with a point forecast. Gaussian Processes (GPs, Rasmussen_Williams_GPML) model the latent function in a non-parametric way, thus avoiding strong assumptions. Furthermore, a Gaussian Process is a Bayesian method: parameters are treated as random variables, to which a prior distribution is assigned. Uncertainty is then updated based on observed data using Bayes’ theorem, resulting in a posterior distribution. The distribution of the data given the parameters is known as the likelihood (Gelman_Hill_Vehtari_2020, Chap. 9).
@@ A GP provides a prior distribution on the space of functions $f:\mathbb{R}_{+}\rightarrow\mathbb{R}$:
The kernel $k$ is a positive definite function (Rasmussen_Williams_GPML, Chap. 4) and encodes the covariance between the values of $f$ at any two instants in time. We assume $f$ to be a smooth function of time. This assumption is encoded by the Radial Basis Function (RBF) kernel:
@@
By defining a prior in the space of functions, GPs do not need regularly spaced observations, therefore we consider a training set of $T$ couples $(t_{1},y_{1}),\ldots,(t_{T},y_{T})$, where, for each $i$, $t_{i}$ is the time stamp and $y_{i}$ is the observation. We aim to generate forecast for time inputs $T+1,\dots,T+h$, where $h$ is known as forecast horizon. We denote the density functions by $p(\cdot)$. The GP provides a prior on the latent vector $\mathbf{f}_{1:T}:=(f(t_{1}),\ldots,f(t_{T}))^{\top}$ which, given a set of time instants $t_{1},\ldots,t_{T}$, encodes the value of the latent function at those times. The prior is a multivariate Gaussian distribution:
@@
Samples of vectors $\mathbf{f}_{1:T}$ are shown in Fig. 1(left) as continuous functions of $t$. Credible intervals display the spread of the prior distribution: at time $t$, the $q$-level credible interval is the interval that contains a proportion $q$ of values of $f(t)$. A priori, the function has the same mean and variance at any time; hence credible intervals are flat (Fig. 1, right).
@@
The latent function $f$, defined by the GP prior in eq. (1), is a time-varying parameter related to the mean of the likelihood function, specified here as a negative binomial (Sec. 3.1) or a Tweedie distribution (Sec. 3.2). For this reason, we require a positive latent function and we ensure this by passing $f$ through the softplus function, defined as $\mathrm{softplus}(x):=\log(1+e^{x})$. The softplus maps negative values of $x$ to small positive values, while it is close to the identity function for $x>2$. A priori, the credible intervals of $\mathrm{softplus}(f)$ are identical for all time instants, non-negative and asymmetric (Fig. 2, left).
@@
where $f_{i}:=f(t_{i})$, $p_{\mathrm{lik}}(\cdot\mid\mathrm{softplus}(f_{i}),\boldsymbol{\theta}_{\mathrm{lik}})$ is either the negative binomial or the Tweedie distribution, and $\boldsymbol{\theta}_{\mathrm{lik}}$ are their hyper-parameters, which we discuss in Secs. 3.1 and 3.2.
@@
The posterior in eq. (4) is available in analytical form only for the Gaussian likelihood. Since our likelihoods are non-Gaussian, we approximate $p(\mathbf{f}_{1:T}~\mid~\mathbf{y}_{1:T})$ with a variational inducing point approximation (Hensman_Fusi_Lawrence_2013) constituted by a Gaussian density, denoted by $q(\mathbf{f}_{1:T}\mid\mathbf{y}_{1:T})$, with tunable mean and covariance parameters. We follow Hensman_Matthews_Ghahramani_2015 to optimize the model parameters as described in A. In what follows we denote the fitted variational approximation by $q(\mathbf{f}_{1:T}\mid\mathbf{y}_{1:T})$.
@@
The distribution of the latent function at future times, shown in eq. (5), is obtained by marginalizing a multivariate Gaussian distribution. Given a likelihood expressed as in eq. (3), $h$-steps ahead forecasts are obtained without autoregressive sampling. We now complete the GP model by specifying the likelihood function.
@@ 3.1 Negative binomial likelihood
The negative binomial is a suitable distribution for intermittent demand (snyder2012forecasting; Kolassa_2016). It addresses the limitations of the Poisson distribution, since it can have both a large probability mass in zero and long tails.
We denote by $\mathrm{NegBin}(y;r,\text{p})$ the density of a negative binomial distribution; $r$ is the number of successes and p denotes, with a slight abuse of notation, the success probability. The mean and the variance of the distribution linearly increase with $r$; p controls the overdispersion. We model $r$ through the GP, by setting $r=\text{softplus}(f)$. Instead, we treat p as a learnable hyper-parameter ($\boldsymbol{\theta}_{\mathrm{lik}}=\{\text{p}\}$) which we keep fixed for all data points of the same time series.
@@
$p_{\mathrm{lik}}\left(y_{i}\mid\mathrm{softplus}(f_{i}),\boldsymbol{\theta}_{\mathrm{lik}}\right)=\mathrm{NegBin}(y_{i};\mathrm{softplus}(f_{i}),\text{p}).$